OpenStudy (anonymous):

Find values of k so that each remainder is 3 the key is -3

OpenStudy (anonymous):

\[(x^3+4x^2+x+k)\div(x+2)\]

OpenStudy (callisto):

By remainder thm, Put x=-2 \[(-2)^{3}+4(-2)^{2}+(-2)+k = 3\] k = 3-6 = -3